OpenStudy (anonymous):

Given function = xcos(x). Need to find the general formula for it's derivative when n=n and test it for n=2 or 3. First five derivatives:- (f)'= cos(x) - x sin(x) (f)''=-2 sin(x) - x cos(x) (f)'''=-3 cos(x) + x sin(x) (f)''''=4 sin(x) + x cos(x) (f)'''''=5 cos(x) - x sin(x) I believe the general formula uses the identity sin (n * π/2), but I've not been able to find the correct general formula which allows me to plug in values of "n" and get the derivatives of xcos(x)

OpenStudy (anonymous):

Any one?

OpenStudy (anonymous):

What I know is that the general formula makes use of sin (n * π/2) & cos (n * π/2)

OpenStudy (anonymous):

y=xcos(x+n(π/2)), y'=?,y''=?,y'''=?,... Is that your question?

OpenStudy (anonymous):

No. y= xcos(x) is the function

OpenStudy (anonymous):

Then what value "n" are you talking about?

OpenStudy (anonymous):

I've evaluated the the first five derivatives to find a pattern. Now I need a general formula for the pattern I've found. A general formula which gives me the derivatives of xcos(x)

OpenStudy (anonymous):

I see. Alright, give me a sec.

OpenStudy (anonymous):

n refers to the number of derivative. For ex. for the first derivative n =1, for the second n=2, etc etc.

OpenStudy (anonymous):

I gots dat.

OpenStudy (anonymous):

For the nth derivative of the function y=xcos(x), for all n∈ℕ,n>0 (yd^n)/(dx^n)=n(sin(x)d^n)/(dx^n)+x(cos(x)d^n)/(dx^n) I think that works...

OpenStudy (anonymous):

u mean n(derivative of sin(x)^n/derivative of x^n) + x( derivative of cos(x)^n)

OpenStudy (anonymous):

Wait a second, let me type this out in LaTeX. XD

OpenStudy (anonymous):

lol thnx man

OpenStudy (anonymous):

Sorry for taking me a while. XD I was testing my knowledge of LaTeX and failing.

OpenStudy (anonymous):

No problem. Lemme check if the equation works. Till then medal awarded ;)

OpenStudy (anonymous):

The formula works.. Thnx alot for your help.